OpenStudy (annebelleschneiders):

Solve the following system of equations and show all work. y = −x^2 + 4 y = 2x + 1

OpenStudy (sweetburger):

substitue the y in the first equation with the y in the second equation you will end up with an equation that looks like 2x+1 = -x^2 +4 then proceed from this

OpenStudy (annebelleschneiders):

what do I do after x^2+x=3/2 ????

OpenStudy (whpalmer4):

Use the quadratic formula? Factor it? But I'm going to save you some trouble and suggest that you've made a mistake somewhere in getting to this point, and should revisit that work first. \[2x+1 = -x^2+4\]I would add \(x^2\) to both sides, subtract \(4\) from both sides, then factor. Set each factor equal to \(0\) and solve for the values of \(x\). Next time, post the math questions in the math section, please..

OpenStudy (annebelleschneiders):

so the answer is x= -3, 1

OpenStudy (whpalmer4):

Let's check them: \[2x+1 = -x^2 + 4\]\[2(-3)+1 = -(-3)^2+4\]\[-6+1 = -9+4\]\[-5=-5\checkmark\] \[2(1) + 1 = -(1)^2+4\]\[2+1 = -1+4\]\[3=3\checkmark\] Good job!